At 11:00a.m. Kate left the Clinton toll booth traveling west at 80 km/h. At 12:00p.m. Juan left the toll booth traveling west at 72 km/h. At what time will the two cars be 98 km apart?
This is what I have so far...Kate: 80(x+1)
Juan: 72x
80(x+1)+72x=98
152x=18
x=.118
Not sure what to do after this..thanks.

Answer by checkley77(12844)

80(x+1)+72x=98
------------------------------------------
Equation should be 80(t+1)-72t=98
80t+80-72t=98
8t=98-80
8t=18
t=18/8
t=2.25 hours they'll be 98 miles apart.
Proof:
80(2.25+1)-72*2.25=98
80*3.25-162=98
260-162=98
98=98
-----------------------------------------
152x=18
x=.118